# Local Lesion Generation is Effective for Capsule Endoscopy Image Data Augmentation in a Limited Data Setting

### v1

@@ 3.2.2 Fine-tuning
The datasets for were constructed from the official training split of the Kvasir Capsule Dataset. Each lesion in the datasets was coupled with a mask based on the available bounding boxes. For each dataset , , the inpainting model was fine-tuned for 500 epochs, incorporating a linear learning rate warm-up over the first 10 epochs. The Adam optimizer was used for both the generator and discriminator, with learning rates set at 0.001 and 0.0001, respectively, and a batch size of 16. Although the learning rate and optimizer settings align with those in [81], the batch size was reduced due to technical constraints. Models were trained on images with a resolution of 256x256, consistent with [81]. The linear learning rate warm-up is a standard practice in fine-tuning deep learning models.
@@ 4.1 Dataset
The dataset is highly imbalanced, as shown in Table 1, which poses a significant challenge for training the classification model [23, 44]. In addition, many lesions are very small and localized, covering only a minor part of the image. Classification of this data is therefore very challenging. Smedsrud et al. [79] achieved just a 25.23% F1-Score macro using DenseNet-161, the best result to date, while Srivastava et al. [80] obtained a 21.78% F1-Score macro with their best model, despite conducting an extensive study of modern classification models on the dataset.
@@ 4.1.1 Data split
4.2 Poisson Blending Data Augmentation
4.2.1 Qualitative results
@@
4.3 Image Inpainting Data Augmentation
The Big LaMa-Fourier model, proposed by Suvorov et al. [81], demonstrates the ability to generate realistic capsule endoscopy images in a zero-shot scenario (Figure 6). It effectively removes lesions, but for IIDA, the reverse behavior is needed. Therefore, in this work, the model is fine-tuned for a reconstruction task, as detailed in Section 3.2.2. The following section will present qualitative results. In GDA, the standard approach to evaluating model performance involves measuring improvements in downstream task metrics compared to a baseline without GDA [52, 109, 38, 15, 86], which will be demonstrated in Section 4.5.
4.3.1 Qualitative results
@@
As demonstrated in Figure 7, the model was provided with identical images, but with varying mask sizes centered in the same image region. The generated samples exhibit significant variability in size, shape, and texture, with lymphangiectasia being particularly well-represented. Additional examples of lymphangiectasia are provided in Figure 8. The model successfully generated plausible lesions for most mask sizes, although it struggled with the smallest mask, failing to produce a lesion when the mask was too small. More examples for each pathology class can be found in Appendix A.
@@ 4.4 Preparation of augmented dataset
For each proposed augmentation method - PBDA and IIDA - three versions of the augmented dataset were created. For each lesion class, synthetic images were generated in sufficient quantities to ensure that the total number of samples per class in the training set reached 2,000, 5,000, or 10,000, depending on the experimental version. If any class exceeded these thresholds, it was limited to 2,000, 5,000, or 10,000 samples accordingly. For simplicity, these datasets are referred to as the 2k, 5k, and 10k datasets in the following sections. This approach ensured that each training set was balanced for the experiments.
@@ 4.5 Classification experiments
To assess our methods’ effectiveness, we conducted experiments on a range of neural network architectures, comparing their performance with standard class imbalance handling techniques. We compared our best-performing model to the previous state-of-the-art approaches on the Kvasir Capsule dataset. Additionally, we investigated the impact of varying augmented sample sizes on classification accuracy, applying our techniques and comparing them with alternative GDA methods. Finally, we evaluated the optimal configurations for our methods to determine their influence on lesion detection performance across the tested methods.
@@ 4.5.2 Study on the effect of augmented sample size
Data Strategy CAFormer-S18 ConvNeXt-Tiny ResNet18 ViT-T/16 AUC F1-score Sens. Spec. AUC F1-score Sens. Spec. AUC F1-score Sens. Spec. AUC F1-score Sens. Spec. Base No strategy Weighting ROS RUS Thresholded Thresholded ROS NVAE 10k 5k 2k LDM 10k 5k 2k PBDA (ours) 10k 5k 2k IIDA (ours) 10k 5k 2k
@@
We conducted additional experiments where synthetic training data was a combination of PBDA-generated data and IIDA-generated data. Figure 12 shows the results for varying proportions of these data generation methods. The optimal performance, particularly in terms of F1-Score macro average, was achieved with a 2k dataset where 20% of the data came from PBDA and 80% from IIDA. Across datasets of 2k, 5k, and 10k samples, there was no clear trend in performance improvement based on the proportion of PBDA to IIDA, except that the 20% PBDA and 80% IIDA configuration usually produced the best results.
@@ 5.2 Image Inpainting Data Augmentation
The IIDA method introduced in this work fine-tunes the LaMa model to convincingly fill missing image regions with realistic lesions. As demonstrated in Figures 7, 8 and 9 IIDA adapts the shape, size, and brightness of generated lesions to blend seamlessly with the surrounding tissue. It also enables the generation of structures with depth, such as lymphangiectasia, which may protrude into the intestinal lumen. Unlike PBDA, IIDA performs well at generating diffuse conditions like bleeding and large objects like foreign bodies.
@@
Results in Section 4.5 show that IIDA consistently outperforms the baseline across all architectures and augmented data sizes. IIDA also consistently achieves higher F1-score than PBDA, albeit with lower sensitivity. Moreover, IIDA shows a smaller reduction in F1-score for the unclear view class compared to PBDA (-6.21 vs. -29.88 p.p., Table 5). These findings, along with the qualitative results in Section 4.3.1 suggest that lesions generated by LaMa are more consistent, but less diverse, which can contribute to overfitting in classification models.
@@ 5.3 Combining PBDA and IIDA
The proposed solutions effectively utilize abundant healthy tissue to enhance lesion recognition. As shown in Table 5, PBDA achieves higher sensitivity than IIDA, though with a lower F1-score. However, Figure 12 and Table 6 demonstrate that a combination of PBDA and IIDA outperforms both methods individually, improving the F1-score to 33.07% and sensitivity to 40.89%. These methods adopt different data augmentation strategies: PBDA applies classical image processing techniques to blend lesions, whereas IIDA generates lesions while conditioned on background tissue. The results suggest that a combination of these methods produces a more diverse dataset, leading to better generalization. The experiments confirm that using both approaches together yields new state-of-the-art augmentation performance.
@@ 5.4 Comparison to other methods
In Sections 4.3.1 and 4.5 we conducted a comprehensive comparison of IIDA and PBDA against alternative methods. Both qualitative and quantitative experiments demonstrated the superiority of our approaches over augmentation using de novo images generated by NVAE and LDM. These generative models occasionally produce global distortions and suffer from data copying issues [11, 58], which is particularly evident when sampling de novo images (Figure 13). PBDA and IIDA may face similar challenges, as PBDA essentially replicates lesions in new locations, and IIDA, as a generative model, might also be prone to lesion copying. However, our methods differ by modifying only selected regions of existing images, limiting the potential for replication errors. The key question is whether our methods can introduce novel lesions into the dataset. As previously discussed, PBDA effectively mitigates dataset bias, while IIDA appropriately adjusts the generated lesions to the surrounding tissue. Although neither method can generate lesions outside the distribution of the training data, the combination of PBDA and IIDA has set a new state-of-the-art, as shown in Table 4, which suggests that the issue of data copying is largely mitigated.
@@ 6 Conclusion
Adrian B. Chłopowiec: Writing - original draft, Validation, Visualization, Software, Methodology, Data curation, Conceptualization. Adam R. Chłopowiec: Writing - original draft, Validation, Software, Methodology, Conceptualization. Krzysztof Galus: Writing - original draft, Validation, Software, Visualization, Data curation. Wojciech Cebula: Writing - original draft, Validation, Software. Martin Tabakov: Writing - original draft, Visualization, Conceptualization, Supervision.
@@
This work was supported by the statutory funds of the Department of Artificial Intelligence, Wroclaw University of Science and Technology.
@@ Acknowledgements
Appendix A Additional samples generated with IIDA

### v2

@@ 3.2.3 Data generation process
However, in this study, to ensure a fair comparison with PBDA, we used samples and ROIs obtained for PBDA using the data preparation procedure outlined in Section 3.1.2. This approach serves as a specific implementation of the general data sampling procedure previously described.
Therefore, the datasets for were constructed from the official training split of the Kvasir Capsule Dataset. For each dataset , , the inpainting model was fine-tuned for 500 epochs, incorporating a linear learning rate warm-up over the first 10 epochs. The Adam optimizer was used for both the generator and discriminator, with learning rates set at 0.001 and 0.0001, respectively, and a batch size of 16. Although the learning rate and optimizer settings align with those in [81], the batch size was reduced due to technical constraints. Models were trained on images with a resolution of 256x256, consistent with [81]. The linear learning rate warm-up is a standard practice in fine-tuning deep learning models.
@@ 4.1 Dataset
The dataset is highly imbalanced, as shown in Table 1, which poses a significant challenge for training the classification model [23, 44]. In addition, many lesions are very small and localized, covering only a minor part of the image. Classification of this data is therefore very challenging. Smedsrud et al [79] achieved just a 25.23% F1-Score macro using DenseNet-161, the best result to date, while Srivastava et al. [80] obtained a 21.78% F1-Score macro with their best model, despite conducting an extensive study of modern classification models on the dataset.
@@ 4.1.1 Data split
4.2 PBDA - qualitative results
@@
More examples for each pathology class can be found in Appendix B.
4.3 IIDA - qualitative results
@@
As demonstrated in Figure 7, the model was provided with identical images, but with varying mask sizes centered in the same image region. The generated samples exhibit significant variability in size, shape, and texture, with lymphangiectasia being particularly well-represented. Additional examples of lymphangiectasia are provided in Figure 8. The model successfully generated plausible lesions for most mask sizes, although it struggled with the smallest mask, failing to produce a lesion when the mask was too small. More examples for each pathology class can be found in Appendix C.
@@ 4.4 Preparation of augmented dataset
In GDA, the standard approach to evaluating model performance involves measuring improvements in downstream task metrics compared to a baseline without GDA [52, 109, 38, 15, 86]. For each proposed augmentation method - PBDA and IIDA - three versions of the augmented dataset were created. For each lesion class, synthetic images were generated in sufficient quantities to ensure that the total number of samples per class in the training set reached 2,000, 5,000, or 10,000, depending on the experimental version. If any class exceeded these thresholds, it was limited to 2,000, 5,000, or 10,000 samples accordingly. For simplicity, these datasets are referred to as the 2k, 5k, and 10k datasets in the following sections. This approach ensured that each training set was balanced for the experiments.
@@ 4.5 Classification experiments
To evaluate the effectiveness of our methods, we conducted experiments using various neural network architectures and compared their performance with standard techniques for addressing class imbalance. This comparison specifically assessed the influence of balancing the training dataset, which is a natural outcome of our proposed data augmentation strategy. We compared our best-performing model to the previous state-of-the-art approaches on the Kvasir Capsule dataset. Additionally, we investigated the impact of varying augmented sample sizes on classification accuracy, applying our techniques and comparing them with alternative GDA methods. Finally, we evaluated the lesion detection performance across the tested methods.
@@ 4.5.2 Study on the effect of augmented sample size
Data Strategy CAFormer-S18 ConvNeXt-Tiny ResNet18 ViT-T/16 AUC F1-score Sens. Spec. AUC F1-score Sens. Spec. AUC F1-score Sens. Spec. AUC F1-score Sens. Spec. Base No strategy Weighting ROS RUS Thresholded Thresholded ROS NVAE 10k 5k 2k LDM 10k 5k 2k PBDA (ours) 10k 5k 2k IIDA (ours) 10k 5k 2k 20% PBDA + 80% IIDA (ours) 10k 5k 2k
@@
We conducted additional experiments where synthetic training data was a combination of PBDA-generated data and IIDA-generated data. Figure 12 shows the results for varying proportions of these data generation methods. The optimal performance, particularly in terms of F1-Score macro average, was achieved with a 2k dataset where 20% of the data came from PBDA and 80% from IIDA. The highest macro average sensitivity of 42.59% was reached using the 5k dataset. Across datasets of 2k, 5k, and 10k samples, there was no clear trend in performance improvement based on the proportion of PBDA to IIDA, except that the 20% PBDA and 80% IIDA configuration usually produced the best results. Similar to IIDA experiments, the 2k dataset yields the best results.
@@ 5.2 Image Inpainting Data Augmentation
The IIDA method introduced in this work fine-tunes an image inpainting model to convincingly fill missing image regions with realistic lesions. As demonstrated in Figures 7, 8 and 9 IIDA adapts the shape, size, and brightness of generated lesions to blend seamlessly with the surrounding tissue. It also enables the generation of structures with depth, such as lymphangiectasia, which may protrude into the intestinal lumen. Unlike PBDA, IIDA performs well at generating diffuse conditions like bleeding and large objects like foreign bodies.
@@
Results in Section 4.5 show that IIDA consistently outperforms the baseline across all architectures and augmented data sizes. IIDA also consistently achieves higher F1-score than PBDA, albeit with lower sensitivity. Moreover, IIDA shows a smaller reduction in F1-score for the unclear view class compared to PBDA (-6.21 vs. -29.88 p.p., Table 5). These findings, along with the qualitative results in Section 4.3 suggest that lesions generated by fine-tuned LaMa are more consistent, but less diverse, which can contribute to overfitting in classification models.
@@ 5.3 Combining PBDA and IIDA
The proposed solutions effectively utilize abundant healthy tissue to enhance lesion recognition. As shown in Table 5, PBDA achieves higher sensitivity than IIDA, though with a lower F1-score. However, Figure 12 and Table 6 demonstrate that a combination of PBDA and IIDA outperforms both methods individually, improving the F1-score to 33.07% using the 2k dataset and sensitivity to 42.59% when training with the 5k dataset. These methods adopt different data augmentation strategies: PBDA applies classical image processing techniques to blend lesions, whereas IIDA generates lesions while conditioned on background tissue. The results suggest that a combination of these methods produces a more diverse dataset, leading to better generalization. The experiments confirm that using both approaches together yields new state-of-the-art augmentation performance.
@@ 5.4 Comparison to other methods
In Sections 4.3 and 4.5 we conducted a comprehensive comparison of IIDA and PBDA against alternative methods. Both qualitative and quantitative experiments demonstrated the superiority of our approaches over augmentation using de novo images generated by NVAE and LDM. These generative models occasionally produce global distortions and suffer from data copying issues [11, 58], which is particularly evident when sampling de novo images (Figure 13). PBDA and IIDA may face similar challenges, as PBDA essentially replicates lesions in new locations, and IIDA, as a generative model, might also be prone to lesion copying. However, our methods differ by modifying only selected regions of existing images, limiting the potential for replication errors. The key question is whether our methods can introduce novel lesions into the dataset. As previously discussed, PBDA effectively mitigates dataset bias, while IIDA appropriately adjusts the generated lesions to the surrounding tissue. Although neither method can generate lesions outside the distribution of the training data, the combination of PBDA and IIDA has set a new state-of-the-art, as shown in Table 4, which suggests that the issue of data copying is largely mitigated.
@@ 6 Conclusion
Adrian B. Chłopowiec: Writing - original draft, Validation, Visualization, Software, Methodology, Data curation, Conceptualization, Investigation. Adam R. Chłopowiec: Writing - original draft, Validation, Software, Methodology, Conceptualization, Formal analysis, Investigation. Krzysztof Galus: Writing - original draft, Validation, Software, Visualization, Data curation. Wojciech Cebula: Writing - original draft, Validation, Software. Martin Tabakov: Writing - original draft, Visualization, Conceptualization, Supervision, Formal analysis.
@@
This work was partially supported by the statutory funds of the Department of Artificial Intelligence, Wroclaw University of Science and Technology and partially supported by the Polish Ministry of Science and Higher Education, NCBR grant No POIR.01.01.01-00-2504/20.
@@ Acknowledgements
Appendix A Algorithms details
Appendix B Additional samples generated with PBDA
Appendix C Additional samples generated with IIDA